How do you solve 4a+3b=-5 and a-b=-3

Answer by mananth(16946)

4 a + 3 b = -5 .............1
1 a -1 b = -3 .............2

multiply (1)by 1
Multiply (2) by 3
4 a + 3 b = -5
3 a -3 b = -9
Add the two equations
7 a = -14
/ 7
a =-2
plug value of a in (1)
4 a + 3 b = -5
-8 + 3 b = -5
3 b = -5 + 8
3 b = 3
b = 1